Shift Summary, 10/10/99 Evening Shift (fwd)

ugur@cuhep.Colorado.EDU
Sun, 10 Oct 1999 22:56:21 -0600 (MDT)

Shift Summary for Shift, 10/10/99vening

On Shift: A. Tripathi, U. Erkarslan

Hours of beam available at all intensities: 0:00
Hours of beam available at nominal intensity: 0:00
Hours of E832 data (special runs not included): 0:00
Hours of special runs (laser scans, muon runs...): 0:00

Accounting of lost time: TeV vacuum valve requires
rewarming of C1 cryo string.
No beam until monday morning

========================================================================

Accesses during shift: none

========================================================================

Overview of shift: No beam, Cosmic run

========================================================================

Description of Shift: 15:55 One of the memories on plane 4 doesn't have
the orange LED blinking.May be the LED
is bad.Try doing a Drt_fresh_start. It
didn't fix the problem


20:35 Ktevana error.

22:30 BobT, Arun, Mike and RickK
are studying for l3
accept fraction problems.Look log book
XXII 72-73 for detail

========================================================================
Items to Follow up on Next ShiftDescription of Shift: